### v1

@@ 6.2.3 Collaborative methods
Li and Xu [114] proposed a novel Bootstrap Knowledge Distillation (BKD) method, which seeks to improve the label qualities gradually, therefore degrade the noise level of the whole dataset. The idea is to transform the distribution of noisy labels to be closer to the distribution of real labels without estimating a noise transition matrix. They applied their method for lung disease classification on CXR images using the CheXpert dataset and the Chest X-ray14 dataset. For the CheXpert dataset, they used the certain part of data to train the teacher model. At the next steps, they employed various strategies to train it on the entire dataset. For handling uncertain labels, they adopted three different strategies they mapped all uncertain labels to 0 (U-Zeros), to 1 (U-Ones), or to the output probability of an auxiliary model (U-SelfTrained). They compared their method with a baseline method which trains a CNN and takes the ensemble of 30 checkpoint using the mean of the output probabilities. They showed that their method which is based on self-train strategy outperformed the baseline method. In particular, despite the fact that the U-SelfTrained strategy seems to be outperforming the other strategies particularly for the “Cardiomegaly” and “Consolidation” detection, the U-ones strategy showed the best results for atelectasis detection, indicating that most of uncertain labels are likely to be positive. For Chest X-ray14 dataset, they used a pre-trained model to select the clean subset. They have chosen the samples whose output probability is larger than 0.55 and ground truth is 1 as certain positive and those whose output probability is smaller than 0.45 and ground truth label is 0 as certain negative while the rest was considered as uncertain labels. When comparing the results of their approach to 3 different published methods, they showed that their method outperformed these baseline methods on most pathologies.
@@ Discussion
One of the ultimate goals of DL models in healthcare is to achieve good generalization performances for wider deployment. This desideratum is of critical importance for the DL models to be employed in real world. However, domain shift is almost inevitable in the medical field. This is mainly because medical data is heterogeneous, exhibiting significant variability due to diverse imaging modalities, patients demographics, and disease characteristics. These factors cause what is call covariate shift. Besides, the data is typically collected in diverse scenarios (e.g., mass screening, city consultations, hospital appointments, etc.), possibly in different countries, implying different annotation guidelines, levels of expertise, etc. These factors cause what we call concept shift. For these reasons, we suspect domain shifts are particularly strong in the medical domain, compared to general-purpose computer vision tasks, where imaging devices (typically cameras) are more homogeneous and concepts (animal species, building types, etc.) are more universal. Facing this domain shift, it is crucial to ensure that DL will perform robustly, reliably and fairly when making predictions about data different from the training data. In this paper, we have presented state-of-the-art strategies for the development of generalized method for medical image classification. Depending on the type of shift, two main categories of methods were identified: covariate shift-based methods and concept shift-based methods.
@@
Hereafter, we will first draw conclusions on the benefits of current DG methods, through the analysis of their results on challenge data (Section 7.1). We will then examine the recent trends in DG development (Section 7.2). Furthermore, we will discuss connections with other research fields (federated learning, fair AI, causal AI, etc.) (Section 7.3). Afterward, we give practical issues to implement and evaluate a DG method (libraries, evaluation strategies, etc.) (Section 7.4). Finally, we end with future directions and promises (subpopulation shift, open DG, etc.) (Section 7.5).
7.1 Lessons learnt from challenges
@@
In histopathology, 17 methods were submitted for the MIDOG challenge [12] final test. These methods were compared to a reference DG approach [209] which reduces covariate shift in the feature space by using adversarial training, belonging to "Representation learning– Adversarial category" in our taxonomy (detailed description of this approach can be found in Section 6.1.2). In addition, a non-DG baseline was considered, named CNN baseline, with the same network topology as the reference approach but only trained using standard image augmentation. Among the submitted methods, three methods [39, 103, 123] were described in this survey and belong to the "Data manipulation–Data augmentation" category (Section 6.1.1).
@@
The findings from the MIDOG 2021 competition suggest that through the use of effective augmentation techniques and sophisticated DL architecture models, domain shift between different whole slide imaging scanners can be addressed to some extent. Despite promising overall performances, the results on unseen Scanners D and F were considerably weaker, indicating that domain shift is not completely covered by the algorithms. This highlights that the problem of DG is not solved yet: there is a need for developing more robust algorithms.
@@
Regarding the recent challenge UBC-OCEAN [^31] https://www.kaggle.com/competitions/UBC-OCEAN which aims to classify ovarian cancer subtypes based on histopathology images, Owkin’s team has won the competition [^32] https://www.kaggle.com/code/jbschiratti/winning-submission. Their solution consisted of using Phikon, Owkin’s foundation model for digital pathology. It is a self-supervised foundation model [56], which consists of a ViT-Base pre-trained with iBOT on 40 million tiles from the TCGA dataset. Specifically, they trained an ensemble of Chowder [43] models (multiple instance learning models) on top of Phikon tile embeddings. Then, they used high entropy predictions to detect outliers. These results suggest that the development of foundation models in the medical field pave the way for improving the generalizability of DL models. In particular, the pre-training strategies are important for enhancing the performances of DL models.
7.2 Trends in DG
@@
For methods dealing with concept shift, we also noted that the number of papers employing learning methods is increasing through the years. We note that some methods use two categories simultaneously (i.e., "data manipulation" and "representation learning", "data adjustment and transformation" and "learning methods").This suggests that combining methods could be also studied in future research to enhance the results.
@@
7.3 Related research to DG
@@ 7.3.3 Causality
7.4 Implementation in generalization research
@@ 7.4.3 Evaluation
7.4.3.1 F1 score
F1 score was used as a metric for evaluating mitotic figure detections in the MIDOG challenge [12]. Overall $F_{1}$ is computed as follow using the counting of all True Postives (TP), False Positives (FP) and False Negatives (FN) detections on slide $i$ for all the $k$ processed slides:
@@ 7.4.3.1 F1 score
7.4.3.2 Quadratic Cohen’s Kappa
This metric [41] compares the performance of the algorithms with the reference standard. It reflects the degree of disagreement, in such a way that more emphasis is given to bigger differences among ratings than to minor differences [180]. It is suitable for multi-category ordinal classification. It was used to assess the algorithms in the PANDA challenge [27].
7.4.3.3 Balanced accuracy
It is defined as the average of recall obtained on each class. This metric was used for assessing the algorithms performances in the ISIC challenge.
7.4.3.4 A-distance
A-distance measures the distribution discrepancy[21]. The smaller the A-distance, the more domain-invariant the features are. Therefore, it is an indicator of how efficient a method is to reduce cross-domain divergence [34].
7.4.3.5 Representation shift
The representation shift (R) is used to quantify the statistical difference between the datasets in the evaluation of DG methods [18, 184]. It computes the differences in the distribution of layer activations of a model between datasets from two domains, capturing the model perceived similarity between the two datasets. The distributions between the two dataset are likely to be similar (small distances) if the model had learnt domain-invariant features.
7.5 Future promises
In this section, we discuss future possibilities for generalization research in medical imaging. We include perspectives for exploring important research area related to DG including datasets, subpopulation shift, open DG, continual DG and unified benchmarking.

### v3

@@ 5.2.3 Collaborative methods
Li and Xu [103] proposed a novel Bootstrap Knowledge Distillation (BKD) method to gradually improve label quality and reduce noise. The method was applied for lung disease classification using the CheXpert and Chest X-ray14 datasets. For the CheXpert dataset, the teacher model was trained on certain data. Then, various strategies were employed to train it on the entire dataset. For handling uncertain labels, three different strategies were adopted: mapping all uncertain labels to 0, to 1, or to the output probability of an auxiliary model. The third strategy outperformed a baseline CNN with an ensemble of 30 checkpoints. For Chest X-ray14 dataset, they used a pre-trained model to select a clean subset. Samples with an output probability larger than 0.55 and a ground truth 1 were considered certain positive, while those with an output probability smaller than 0.45 and ground truth label 0 were certain negatives; the rest were uncertain labels. Their method showed good performances, outperforming state-of-the-art methods on most pathologies.
Public medical datasets for generalization research
There exist many public datasets which have been adopted for generalization research in the medical field. For example, to prevent domain shifting, some mutlti-institutional datasets have been proposed for segmentation problems [161, 162] and image reconstruction [163]  [^32] https://brain-development.org/ixi-dataset/,https://www.kaggle.com/c/second-annual-data-science-bowl/data,http://mridata.org/. In this section, we present the publicly available datasets that were used for classification experiments in the selected articles. Table 2 summarizes these public medical datasets which can be used for generalization research. For a more comprehensive understanding for readers, we will give brief details for public datasets available as part of a challenge.
MIDOG datasets Mitosis Domain Generalization (MIDOG) dataset targets the detection of mitotic figures in histopathology images under domain shift regime.
  • MIDOG 2021 dataset [164]: This dataset was part of the MICCAI MIDOG 2021 challenge which aims to evaluate methods that mitigate domain shift and derive scanner-agnostic algorithms. It addresses DG in histopathology. The main task was mitosis detection in breast cancer. The challenge dataset features 300 cases, 6 scanners, and more than 2500 mitosis. The domains are defined by scanner types.
CAMELYON datasets Cancer Metastases in Lymph nodes challenge (CAMELYON) datasets target the automated detection of cancer metastases in Whole-Slide Images (WSIs) of sentinel lymph nodes.
LUNA-16: The goal of LUNA-16 challenge is the automated detection of pulmonary nodules in thoracic Computed Tomography (CT) scans [168]. This challenge use data from a large public LIDC-IDRI dataset  [150]. More precisely, scans with a slice thickness greater than 2.5 mm were excluded. The resulting dataset contains 888 CT scans.
PANDA dataset: The goal of Prostate Cancer Grade Assessment (PANDA) challenge [169] is the diagnosis of prostate cancer in biopsies. It aims to develop AI algorithms for Gleanson grading. In total, the PANDA dataset comprises 12,625 WSIs of prostate biopsies retrospectively collected from 6 different sites for algorithm development, tuning and independent validation. Cases for development, tuning and internal validation originated from two European (EU) centers: Radboud University Medical Center, Nijmegen, the Netherlands and Karolinska Institutet, Stockholm, Sweden. The external validation data consisted of a US (741 cases) and an EU set (330 cases).
Dataset Modality (Organs) Number of cases Reference MIDOG 2021 dataset [164] Histopathology (Breast) 300 images https://midog2021.grand-challenge.org/ VGH [170] 5,920 images https://tma.im/tma_portal/C-Path/supp.html NKI [170] 8,337 images Camelyon16 WILDS dataset [165] 399 WSI https://camelyon16.grand-challenge.org/Data/ PatchCamelyon [166] 327,680 color images https://patchcamelyon.grand-challenge.org/ Camelyon17 WILDS dataset [167] 1000 WSI https://camelyon17.grand-challenge.org/Data/ TCGA-BRCA 1,098 cases https://portal.gdc.cancer.gov/projects/TCGA-BRCA BACH dataset Microscopy,Histopathology (Breast) 400 microscopy images 30 WSI https://iciar2018-challenge.grand-challenge.org/Dataset/ UBC-OCEAN Histopathology (Ovaries) 538 training images https://www.kaggle.com/competitions/UBC-OCEAN/data PANDA Histopathology (Prostate) 12,625 WSIs https://www.kaggle.com/c/prostate-cancer-grade-assessment DiagSet-B 4675 scans https://github.com/michalkoziarski/DiagSet SICAPv2 [171] 155 biopsies (95 patients) https://data.mendeley.com/datasets/9xxm58dvs3/1 TUPAC-16[172] Histopathology (Colon) 1076 cases https://tupac.grand-challenge.org/Dataset/ Kather16 [173] Histopathology (Colon) 5,000 patches https://zenodo.org/records/53169 Kather19 [174] 100,000 patches http://dx.doi.org/10.5281/zenodo.1214456 CRC-TP [175] 196,000 patches https://warwick.ac.uk/TIAlab/data/crchistolabelednucleihe/. IHC [176] 1,376 images http://fimm.webmicroscope.net/supplements/epistroma CRC-VALHE-7K 7,180 image patches https://zenodo.org/records/1214456 Stanford-CRC [177] 66,578 image tiles https://github.com/rikiyay/MSINet CRC-DX-TRAIN dataset 93,408 image tiles https://zenodo.org/records/2530835#.XwCkDZNKhTY CRC-DX-TEST dataset 99,904 image tiles https://zenodo.org/records/2530835#.XwCkDZNKhTY Chaoyang Dataset [80] 6,160 WSI https://bupt-ai-cz.github.io/HSA-NRL/ DigestPath2019 [178] 690 patients https://digestpath2019.grand-challenge.org/ TCGA-LGG Histopathology (Brain) 516 cases https://portal.gdc.cancer.gov/projects/TCGA-LGG TCGA-GBM 617 cases https://portal.gdc.cancer.gov/projects/TCGA-LGG TCGA-LUAD Histopathology (Lung) 585 cases https://portal.gdc.cancer.gov/projects/TCGA-LUAD TCGA-LUSC 504 cases https://portal.gdc.cancer.gov/projects/TCGA-LUSC ADNI-1 [122] MRIs (Brain) 748 subjects https://adni.loni.usc.edu/ ADNI-2 [122] 708 subjects https://adni.loni.usc.edu/ AIBL [123] 549 subjects https://adni.loni.usc.edu/ ISIC-2017 [179] Dermoscopic images 2,750 images https://challenge.isic-archive.com/data/ ISIC-2019[180] 33,569 images HAM10000 (HAM)[180] 10,000 images https://dataverse.harvard.edu/dataset.xhtml?persistentId=doi:10.7910/DVN/DBW86T Dermofit (DMF) 1,300images https://homepages.inf.ed.ac.uk/rbf/DERMOFIT/datasets.htm Derm7pt (D7P) [107] 2,000 images http://derm.cs.sfu.ca INbreast [181] Mammography (Breast) 115 cases https://www.kaggle.com/datasets/ramanathansp20/inbreast-dataset OPTIMAM dataset [182] 179,326 cases https://medphys.royalsurrey.nhs.uk/omidb/ BCDR [183] 3,703 digitised film mammograms https://www.medicmind.tech/cancer-imaging-data LIDC-IDRI [150] CT(Lung) 1,018 scans (1,010 subjects) https://wiki.cancerimagingarchive.net/pages/viewpage.action?pageId=1966254 LUNA-16 888 scans https://luna16.grand-challenge.org/Data/ LUNA-DG [30] 887 scans https://github.com/meisun1207/LUNA-DG NLST [184, 185] 25,681 patients (77,040 images) https://cdas.cancer.gov/nlst/ COVID19-Diag 226 CT volumes https://github.com/MLMIP/COVID19-Diag ChestX-ray14 (NIH) [127] X-ray (Chest) 112,120 images (30,805 subjects) https://nihcc.app.box.com/v/ChestXray-NIHCC CheXpert (CXP) [128] 224,316 images (65,240 subjects) https://stanfordmlgroup.github.io/competitions/chexpert/ MIMIC-CXR (MMC) [126] 377,110 images (65,179 subjects) https://physionet.org/content/mimic-cxr/2.0.0/ PadChest [129] 160,000 images(67,000 subjects) https://bimcv.cipf.es/bimcv-projects/padchest/ Open-i dataset [186] 8,121 images http://openi.nlm.nih.gov/ Tawsifur [187, 188] 931 images https://www.kaggle.com/datasets/tawsifurrahman/covid19-radiography-database Skytells 1,017 images https://github.com/skytells-research/COVID-19-XRay-Dataset UK Biobank retinal photography dataset [189] Fundus photography (Eye) 58,700 patients https://www.ukbiobank.ac.uk/ EyePACs 88,702 images https://www.kaggle.com/c/diabetic-retinopathy-detection/data APTOS 3,662 images https://www.kaggle.com/c/aptos2019-blindness-detection Messidor 1,200 images https://www.adcis.net/en/third-party/messidor/ PALM 1,200 images https://palm.grand-challenge.org/ AV-DRIVE [190] 40 Images https://drive.grand-challenge.org/ LES-AV 22 images https://figshare.com/articles/dataset/LES-AV_dataset/11857698 HRF 45 images https://www5.cs.fau.de/research/data/fundus-images/ REFUGE 1,200 images https://refuge.grand-challenge.org/ REFUGE2 2,000 images https://refuge.grand-challenge.org/ STARE 20 images https://cecas.clemson.edu/~ahoover/stare/probing/index.html RIGA 750 images https://academictorrents.com/details/eb9dd9216a1c9a622250ad70a400204e7531196d DDR 13,673 images https://drive.google.com/drive/folders/1z6tSFmxW_aNayUqVxx6h6bY4kwGzUTEC RIMONEv2 [191] 455 images https://medimrg.webs.ull.es/ FGADR 1,842 images https://csyizhou.github.io/FGADR/ Endovis Challenge dataset [192] Endoscopic (Abdominal organs) 8 sequences https://endovissub2017-roboticinstrumentsegmentation.grand-challenge.org/ Heidelberg colorectal dataset [193] Laparoscopy (Colorectum) 30 laparoscopic videos https://robustmis2019.grand-challenge.org/Data/ CholecSeg8k [194] Laparoscopy (Abdomen) 17 videos from Cholec80 dataset http://camma.u-strasbg.fr/datasets SurgicalActions [195] Laparoscopy (Gynecologic organs) 160 videos http://ftp.itec.aau.at/datasets/SurgicalActions160/ Cataract-101 [196] Video (Eye) 101 cataract surgeries http://ftp.itec.aau.at/datasets/ovid/cat-101/ PathVQA [197] Multiple modalities (multi-organs) 4,998 pathology images (multi-organs) https://github.com/UCSD-AI4H/PathVQA VQA-RAD[198] Radiology (multi-organs) 315 radiology images. https://osf.io/89kps/ OrganCMNIST CT (Abdomen) 23,583 https://medmnist.com/
Table 2: Public medical datasets used for generalization research. All hyperlinks in this paper were retrieved on 22 March 2024.
@@ Discussion
One of the ultimate goals of DL models in healthcare is to achieve good generalization performances for wider deployment. This desideratum is of critical importance for DL models to be employed in the real world. However, domain shift is almost inevitable in the medical field. Medical data is heterogeneous, exhibiting significant variability due to diverse imaging modalities, patients demographics, and disease characteristics. These factors are responsible for the occurence of covariate shift. Besides, data is typically collected in diverse scenarios (e.g., mass screening, city consultations, hospital appointments, etc.), possibly in different countries, implying different annotation guidelines, levels of expertise, etc. These factors lead to the manifestation of concept shift. For these reasons, we suspect domain shifts are particularly pronounced in the medical domain, compared to general-purpose computer vision tasks, where imaging devices (typically cameras) are more homogeneous and concepts (animal species, building types, etc.) are more universal. Facing this domain shift, it is crucial to ensure that DL will perform robustly, reliably and fairly when making predictions about data different from the training data. In this paper, we have presented state-of-the-art strategies for the development of generalized method for medical image classification. Depending on the type of shift, two main categories of methods were identified: covariate shift-based methods and concept shift-based methods.
@@
Hereafter, we will first present our analysis of the methods and examine the recent trends in DG development (Section 7.1). Next, we will discuss connections with other research fields (federated learning, fair AI, causal AI, etc.) (Section 7.2). Afterward, we will address practical issues in implementing and evaluating a DG method (libraries, evaluation strategies, etc.) (Section 7.3). Finally, we will conclude with future directions and promises (subpopulation shift, open DG, etc.) (Section 7.4), and limitations (Section 7.5).
7.1 What are the state-of-the-art methods in medical image classification for generalization research in the literature?
Our present taxonomy is based on the analysis of existing studies in generalization research in medical images. For covariate shift, we identified three main categories: data manipulation, representation learning, and learning methods. For concept shift, our taxonomy proposes three main categories: data adjustment and transformation, learning methods, and collaborative methods. In this section, we will first start with a critical analysis of the current state of the field and we will draw conclusions on the benefits of current DG methods, through the analysis of their results on challenge data (Section 7.1.1). We will then examine the recent trends in DG development (Section 7.1.2).
7.1.1 Lessons learnt from challenges
Our survey has shown that the generalization problem is common for a variety of modalities: histopathology, X-ray, fundus photographs, ultrasound, etc.
The authors of the reviewed studies used different protocols and datasets splits, making it difficult to compare the results effectively. For a comprehensive comparison, we have included the results of the methods proposed in the MIDOG challenge (Table 3).
Article Category Subcategory F1 score on the preliminary test set F1 score on the final test set Other test set Almahfouz Nasser et al. [29] Data Manipulation Data homogenization 0.0030 Li et al. [19] Data augmentation Average accuracy 0.6285 Chung et al. [40] 0.7548 0.7243 Dexl et al. [37] 0.7138 0.6963 Lafarge and Koelzer [36] 0.6828 0.6319 Long et al. [38] 0.7500 0.7010 Wilm et al. [45] Representation learning Adversarial 0.750 0.7183 Razavi et al. [69] Learning strategies Multi-task learning 0.7492 0.7064
Table 3: Generalizing Methods proposed for the MIDOG challenge.
@@ 7.1.1 Lessons learnt from challenges
A total of 17 methods were submitted for the MIDOG challenge [164] final test. These methods were compared to a reference DG approach [45], presented in Section 5.1.2 (penultimate row in Table3). This approach reduces covariate shift in the feature space by using adversarial training, belonging to “Representation learning– Adversarial category” in our taxonomy. In addition, a non-DG baseline was considered, named CNN baseline, with the same network topology as the reference approach but only trained using standard image augmentation. Among the submitted methods to the final phase, four methods [40, 37, 36, 38] were described in this survey and belong to the “Data manipulation–Data augmentation” category (Section 5.1.1), and one method Razavi et al. [69] belongs to the “Learning strategies– Multi-task learning” category.
In contrast to the aforementioned methods, Li et al. [19] considered single DG setting using the MIDOG dataset. They trained their model using data coming from one scanner domain and then evaluated it on all the other domains. This process was repeated with each domain serving as the training data. Finally, they have computed the mean performance across all unseen domains.
@@
The findings from the MIDOG 2021 competition suggest that through the use of effective augmentation techniques and sophisticated DL architecture models, domain shift between different whole slide imaging scanners can be addressed to some extent. Despite promising overall performances, the results on unseen Scanners were considerably weaker, indicating that domain shift is not completely covered by the algorithms. This highlights that the problem of DG is not solved yet: there is a need for developing more robust algorithms.
@@
In the same context of histopathology, a more recent challenge UBC-OCEAN [^33] https://www.kaggle.com/competitions/UBC-OCEAN aimed to classify ovarian cancer subtypes based on histopathology images. Owkin’s team has won the competition [^34] https://www.kaggle.com/code/jbschiratti/winning-submission. Their solution consisted of using Phikon, Owkin’s foundation model for digital pathology. It is a self-supervised foundation model [199], which consists of a ViT-Base pre-trained with iBOT on 40 million tiles from the TCGA dataset. Specifically, they trained an ensemble of Chowder [200] models (multiple instance learning models) on top of Phikon tile embeddings. Then, they used high entropy predictions to detect outliers. These results suggest that the development of foundation models in the medical field pave the way for improving the generalizability of DL models. In particular, the pre-training strategies are important for enhancing the performances of DL models.
7.1.2 Trends in DG
@@
For methods dealing with concept shift, we also noted that the number of papers employing learning methods is increasing through the years. We note that some methods use two categories simultaneously (i.e., “data manipulation” and “representation learning”, “data adjustment and transformation” and “learning methods”).This suggests that combining methods could be also studied in future research to enhance the results.
@@
7.2 What are the related areas the generalization research?
@@ 7.2.3 Causality
7.3 What are the best practices for implementing generalization techniques in research?
@@ 7.3.3 Evaluation
F1 score F1 score was used as a metric for evaluating mitotic figure detections in the MIDOG challenge [164]. Overall $F_{1}$ is computed as follow using the counting of all True Postives (TP), False Positives (FP) and False Negatives (FN) detections on slide $i$ for all the $k$ processed slides:
@@
Quadratic Cohen’s Kappa This metric [216] compares the performance of the algorithms with the reference standard. It reflects the degree of disagreement, in such a way that more emphasis is given to bigger differences among ratings than to minor differences [217]. It is suitable for multi-category ordinal classification. It was used to assess the algorithms in the PANDA challenge [169].
Balanced accuracy It is defined as the average of recall obtained on each class. This metric was used for assessing the algorithms performances in the ISIC challenge.
A-distance A-distance measures the distribution discrepancy[218]. The smaller the A-distance, the more domain-invariant the features are. Therefore, it is an indicator of how efficient a method is to reduce cross-domain divergence [47].
Representation shift The representation shift (R) is used to quantify the statistical difference between the datasets in the evaluation of DG methods [64, 10]. It computes the differences in the distribution of layer activations of a model between datasets from two domains, capturing the model perceived similarity between the two datasets. The distributions between the two dataset are likely to be similar (small distances) if the model had learnt domain-invariant features.
7.4 What are the key challenges and future promises for generalization research?
In this section, we discuss future possibilities for generalization research in medical imaging. We include perspectives for exploring important research area related to DG including datasets, modelling pipeline strategies, subpopulation shift, open DG, continual DG, unified benchmarking, privacy concerns and multimodal DG.